# FlightBERT++: A Non-autoregressive Multi-Horizon Flight Trajectory Prediction Framework

### v2

@@ Introduction
The DPD receives two input vectors, i.e., the differential embeddings $\mathbf{D}_{t-k+2:t}$ and multi-horizon context representations $\mathbf{C}_{t+1:t+n}$. Specifically, the differential embeddings $\mathbf{D}_{t-k+2:t}$ of the $O_{t-k+1:t}$ are extracted by a Conv1D neural network, which severs as the prompting to learn the transition patterns of the differential sequence. Then, the extracted $\mathbf{D}_{t-k+2:t}$ and multi-horizon context representations $\mathbf{C}_{t+1:t+n}$ are jointly fed into the differential-prompted decoder to generate the final outputs as Eq. (5), where the $\mathbf{P}_{t+1:t+n}^{BE}$ is the BE representation of the predicted differential sequence of future trajectory. Finally, the $\mathbf{P}_{t+1:t+n}^{BE}$ is transformed into decimals to reconstruct the predicted trajectory based on the $O_{t}$.
$\mathbf{P}_{t+1:t+n}^{BE}=\operatorname{DPD}([\mathbf{D}_{t-k+2:t},\mathbf{C}_{t+1:t+n}])$ (5)
Note that the inputs and the outputs of the FlightBERT++ are both the BE representation of the trajectory attributes. Therefore, the optimizing objective of the FlightBERT++ is also formulated as the MBC task. The FlightBERT++ framework is trained using the Binary Cross Entropy (BCE) loss function, as that in MBC tasks. More details of the BE representations can be found in (Guo et al. 2023).
Compared with existing multi-horizon FTP approaches, the proposed framework (i) innovatively design a HACG to generate multi-horizon contexts directly, (ii) employ the Transformer-based architecture to conduct the backbone network of the trajectory encoder and differential-prompted decoder, which are the core ideas in enabling non-autoregressive prediction.
@@
Trajectory Encoder
As illustrated in Figure 1, the trajectory encoder is composed of three modules, including the Conv1D-based Trajectory Point Embedding (TPE) module, Transformer-based temporal modeling (TTM) module, and Attention-based Sequence Aggregation (ASA) module. Specifically, the Conv1D-based TPE module projects the BE representation into high-dimensional embedding space to learn discriminative spatial features of the trajectory points, while the Transformer-based module is employed to capture the temporal correlations of the observation sequence. In succession, the outputs of the Transformer module are further fed into the ASA module to generate the trajectory-level embedding and extract the semantic representation over the whole observation sequences.
To learn the informative spatial features of the trajectory points, a Conv1D-based channel-mix trajectory point embedding module is proposed to project the BE representations of trajectory points into embedding space. As illustrated in Figure 2, for a trajectory point, a joint embedding is obtained by concatenating its attribute-wise BE representations, which is further fed into a Conv1D layer to learn trajectory point embedding. The channel-mix strategy not only retains the global correlations of the trajectory attributes, but also is beneficial to learn the local features among the bits of the joint BE representation.
The TTM module is implemented by the stacked Transformer blocks. Compared with the RNN-based architectures, the Transformer is primarily implemented by the self-attention mechanism, which enables the model to learn the temporal correlations of the observations in a non-autoregressive manner. The ASA module performs a weighted sum operation in the temporal dimension, in which the attention weights of the trajectory points are generated as Figure 2. In this way, the trajectory-level embedding $\mathbf{Traj}_{enc}$ can be obtained from the trajectory encoder, which is expected to capture the temporal-spatial semantic representation of the observations.
Horizon-Aware Context Generator
In general, the autoregressive nature of conventional multi-horizon FTP approaches degrades the efficiency in the training and inference process, especially in high temporal resolution conditions, which may be not suitable for real-time forecasting scenarios. Towards this gap, a horizon-aware context generator is innovatively designed to support multi-horizon predictions in a non-autoregressive manner by generating multi-horizon contexts directly.
@@
The core idea of the HACG is to leverage the trajectory-level embedding $\mathbf{Traj}_{enc}$ and horizon embeddings $\mathbf{He}_{1:n}$ to generate informative multi-horizon context representations. Specifically, the $\mathbf{Traj}_{enc}$ can be regarded as a high-dimensional representation that implies the global features of the observations, while the horizon embeddings $\mathbf{He}_{1:n}$ provide the different semantic representations for each predicted horizon. Based on this assumption, the concatenation operation and MLP block are employed to fuse these vectors and generate the multi-horizon context representations. By integrating the prior horizon knowledge, the HACG can be aware of different horizons and generate the multi-horizon context representations directly via only one-pass inference.
Differential-prompted Decoder
To mitigate the high-bits prediction errors of the BE representation, the differential prediction paradigm is introduced into the FlightBERT++ framework, i.e., the objective of the decoder is to predict the differential values instead of the raw absolute values. However, it is challenging to learn the transition patterns of differential sequence from the observations sequence because the differential operation may ignore some geographical and kinematical features of the trajectory attributes. To this end, as illustrated in Figure 1, a DPD is proposed to reduce the learning difficulty of the network by integrating the differential prompted mechanism.
@@
Specifically, the DPD consists of two modules, i.e., masked-Transformer, and predictor. Firstly, the differential sequence is calculated from the observed sequence and encoded into BE representations which can be consistent with the form of the outputs. Similar to the TPE module, these BE representations are further fed into a Conv1D-based embedding layer to learn the high-dimensional differential embeddings $\mathbf{D}_{t-k+2:t}$. Secondly, the $\mathbf{D}_{t-k+2:t}$ is concatenated with the multi-horizon context representations $\mathbf{C}_{t+1:t+n}$ along the temporal dimension as a prompt to learn the transition patterns of the predicted differential sequence. In succession, the concatenated vectors further fed into the masked-Transformer module to build the inter- and intra-temporal correlations across observation and the multi-horizon predictions. The architecture of the masked-Transformer module is similar to the TTM module except it employs the masked self-attention mechanism to ensure the temporal specificities of the sequence. Finally, the predictor is composed of a Fully Connected (FC) layer and the Sigmoid activation, which is applied to output BE representations for the multi-horizons.
Experiments
Experimental Settings
Dataset Preprocessing and Description
@@
To validate the proposed framework, a real-world flight trajectory dataset was collected from an ATC system. The dataset contains a total of 9 days of trajectory data with 20-second intervals, in which the range of the interested (ROI) longitude and latitude are $[94.616\degree,113.689\degree]$ and $[19.305\degree,37.275\degree]$, respectively. The key attributes of the flight trajectory are extracted from the raw data to support the experiments, including timestamps, call sign, longitude, latitude, altitude, and velocity in $x,y,z$ directions. After preprocessing, a total of 8643 flight trajectories in the dataset that further split into train, validate, and test subsets. Specifically, the trajectory of the first 7 days is used to train the FTP models while $8^{th}$ and $9^{th}$ are for validation and testing, respectively.
Comparison Baselines
@@
In this work, a total of 5 competitive approaches serve as baselines to validate the effectiveness of the proposed FlightBERT++. Moreover, the baselines are divided into two groups to conduct comparisons by multi-horizon prediction styles. Group A performs iterative multi-horizon prediction, i.e., the model only predicts the result for the next time step that further serves as a pseudo-observation for multi-horizon inference, including LSTM (Shi et al. 2018), Transformer (Vaswani et al. 2017), Kalman-Filter (Kalman 1960), FlightBERT (Guo et al. 2023). Group B is a direct multi-horizon prediction approach, i.e., forecasts the results in a single inference process for multiple time steps, which is implemented by the LSTM+Attention model (Bahdanau, Cho, and Bengio 2015).
Experimental Configuration
Based on ATC work, the quantization precision $0.001\degree$ (about 110 meters) is selected to adapt the transformation between BE representation and the decimals. In this resolution, the input of the proposed framework is a 78-dimensional vector (BE representation) while the output is a 48-dimensional vector. The embedding size of the trajectory points, horizons, and differentials is set to 128, i.e., 128 channels for the Conv1D networks in the proposed framework. The number of the Transformer blocks in the trajectory encoder and differential-prompted decoder are both set to 4. The number of hidden states of the Transformer blocks is set to 768. An attention operator with 4 heads is applied to all Transformer blocks in the proposed framework. In the experiments, we use the latest 3-minute observations to predict the flight status of the future 5 minutes, i.e., predicting 15 trajectory points based on 9 observed trajectory points. The Adam optimizer with an initial learning rate of $10^{-4}$ is applied to train all the above deep learning-based models.
Style Methods Hor. MAE $\downarrow$ MAPE (%) $\downarrow$ RMSE $\downarrow$ MDE $\downarrow$
@@
Direct LSTM+Attention 1 0.0064 0.0068 1.98 0.0059 0.0249 0.30 0.0109 0.0138 5.96 1.09
@@
Table 1: The experimental results of the proposed framework and baselines.
Evaluation Metrics
In this work, the Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE) are applied to evaluate the proposed methods and baselines, which are the common criteria for the FTP tasks. In addition, to intuitively validate the FTP performance in the three-dimensional (3D) airspace, the Mean Distance Error (MDE) is proposed to evaluate the Euclidean distance of the predictions and the actual trajectory points. Specifically, the positional attributes of both predictions and ground truth are projected to the earth-centered and earth-fixed (ECEF) coordinate system, and the distance error is measured to evaluate the model performance as below:
where $p,p^{\prime}$ are the transformed values of ground truth and prediction in the ECEF coordinate system. $\Phi(\cdot)$ is the calculation function of Euclidean distance in the 3D airspace.
Furthermore, the Mean Time Costs (MTC) metric is proposed to evaluate the computational performance, which is defined as follows:
$MTC=\frac{1}{N}\sum_{i=1}^{N}time\_costs_{i}^{h}$ (10)
where $N$ is the number of samples in the test process, $time\_costs_{i}^{h}$ represents the time cost for $h$ prediction horizons of sample $i$. In this phase, the batch size of all evaluation models is set to 1 to ensure comparison fairness.
Result and Discussions
Results and Quantitative Analysis
Overall Performance of FTP

### v4

@@ III-C The Proposed Framework
The DPD receives two input vectors, i.e., the differential embeddings $\mathbf{D}_{t-k+2:t}$ and multi-horizon context representations $\mathbf{C}_{t+1:t+n}$. Specifically, the differential embeddings $\mathbf{D}_{t-k+2:t}$ of the $O_{t-k+1:t}$ are extracted by a Conv1D neural network, which severs as the prompting to learn the transition patterns of the differential sequence. Then, the extracted $\mathbf{D}_{t-k+2:t}$ and multi-horizon context representations $\mathbf{C}_{t+1:t+n}$ are jointly fed into the differential-prompted decoder to generate the final outputs as Eq. (5), where the $\mathbf{P}_{t+1:t+n}^{GC}$ is the GC representation of the predicted differential sequence of future trajectory. Finally, the $\mathbf{P}_{t+1:t+n}^{GC}$ is transformed into decimals to reconstruct the predicted trajectory based on the $O_{t}$.
$\mathbf{P}_{t+1:t+n}^{GC}=\operatorname{DPD}([\mathbf{D}_{t-k+2:t},\mathbf{C}_{t+1:t+n}])$ (5)
Note that the inputs and the outputs of the proposed framework are both the GC representations of the trajectory attributes. Therefore, the optimizing objective of the proposed framework is also formulated as the MBC task. The FlightBERT++ framework is trained using the Binary Cross Entropy (BCE) loss function, as that in conventional MBC tasks. More details of the BE representations and MBC-based FTP task can be found in our previous work [9].
Compared to existing multi-horizon FTP approaches, the proposed FlightBERT++ framework (i) innovatively designed a HACG to generate multi-horizon contexts directly, (ii) employ the Transformer-based architecture to conduct the backbone network of the trajectory encoder and differential-prompted decoder, which are the core ideas in enabling non-autoregressive prediction.
III-D Trajectory Encoder
As illustrated in Figure 2, the trajectory encoder is composed of three modules, including the Conv1D-based Trajectory Point Embedding (TPE) module, Transformer-based temporal modeling module, and Attention-based Sequence Aggregation (ASA) module. Specifically, the Conv1D-based TPE module projects the GC representation into high-dimensional embedding space to learn discriminative spatial features of the trajectory points, while the Transformer-based module is employed to capture the temporal correlations among the observation sequence. In succession, the outputs of the Transformer module are further fed into the ASA module to generate the trajectory-level embedding and extract the semantic representation over the whole observation sequences.
@@ III-D Trajectory Encoder
To capture the informative spatial features of the trajectory points, the linear-based attributes embedding block is designed in [9] to learn attribute embeddings in a channel-independent manner. However, the channel-independent approaches suffer from two limitations. i) Since strong spatial correlations are implied among the trajectory attributes, learning the attributes embedding separately might not fully consider the global features. ii) The parallel linear layers in the attributes embedding block burden the complexity of the model.
To address the aforementioned limitations, in this work, a simple yet efficient Conv1D-based channel-mix trajectory point embedding module is proposed to project the GC representations of trajectory points into high-dimensional embedding space. As illustrated in Figure 3, for a trajectory point, a joint embedding is obtained by concatenating its attribute-wise GC representations, which is further fed into a Conv1D layer to learn trajectory point embedding. On the one hand, the channel-mix strategy retains the global correlations of the trajectory attributes. On the other hand, the Conv1D neural network is beneficial for learning the local features among the bits of the joint GC representation.
The Transformer-based temporal modeling module is implemented by the stacked Transformer blocks. Compared to the RNN-based architectures, the Transformer is primarily implemented by the self-attention mechanism, which enables the model to learn the temporal correlations of the observations in a non-autoregressive manner. The ASA module performs a weighted sum operation in the temporal dimension, in which the attention weights of the trajectory points are generated as Figure 3. In this way, the trajectory-level embedding $\mathbf{Traj}_{enc}$ can be obtained from the trajectory encoder, which is expected to capture the temporal-spatial semantic representation of the observations.
III-E Horizon-Aware Context Generator
Conventional multi-horizon approaches usually perform the prediction in an autoregressive manner, i.e., predicting $p_{t+1}$ based on observation and generating the $p_{t+2}$ according to the $p_{t+1}$ sequentially. However, the autoregressive nature of these approaches degrades the efficiency in the training and inference process, especially in high temporal resolution conditions, which may be not suitable for real-time forecasting scenarios. Towards this gap, a horizon-aware context generator is innovatively designed to support multi-horizon predictions in a non-autoregressive manner by generating multi-horizon contexts directly.
@@ III-E Horizon-Aware Context Generator
The core idea of the HACG is to leverage the trajectory-level embedding $\mathbf{Traj}_{enc}$ and horizon embeddings $\mathbf{He}_{1:n}$ to generate informative multi-horizon context representations. Specifically, the $\mathbf{Traj}_{enc}$ can be regarded as a high-dimensional representation that implies the global features of the observations, while the horizon embeddings $\mathbf{He}_{1:n}$ provide the different semantic representations for each predicted horizon. Based on this assumption, the concatenation operation and MLP block are employed to fuse these vectors and generate the multi-horizon context representations. By integrating the prior horizon knowledge, the HACG is able to be aware of different horizons and generate the multi-horizon context representations directly via only one-pass inference.
III-F Differential-prompted Decoder
Figure 4: The detailed implementation of the Differential-prompted Decoder.
@@ III-F Differential-prompted Decoder
As described in Section III-B, to suppress outliers caused by high-bit misclassifications, the differential prediction paradigm is introduced into the FlightBERT++ framework. In practice, it is challenging to learn the transition patterns of differential sequence from the observations sequence because the differential operation may ignore some geographical and kinematical features of the trajectory attributes. To this end, as illustrated in Figure 2 and Figure 4, a differential-prompted decoder is proposed to reduce the learning difficulty of the network by integrating the differential prompted mechanism. Specifically, the differential-prompted decoder consists of two modules, i.e., masked-Transformer, and predictor. Firstly, the differential sequence is calculated from the observed sequence and encoded into GC representations which can be consistent with the form of the outputs. Similar to the TPE module, these GC representations are further fed into a Conv1D-based embedding layer to learn the high-dimensional differential embeddings $\mathbf{D}_{t-k+2:t}$. Secondly, the learned differential embeddings $\mathbf{D}_{t-k+2:t}$ are concatenated with the multi-horizon context representations $\mathbf{C}_{t+1:t+n}$ along the temporal dimension as a prompt to learn the transition patterns of the differential sequence.
In succession, the concatenated vectors further fed into the masked-Transformer module to build the inter- and intra-temporal correlations across observation and the multi-horizon predictions. The architecture of the masked-Transformer module is similar to the Transformer-based temporal modeling module except it employs the masked self-attention mechanism to ensure the temporal specificities of the sequence. Finally, the predictor is composed of a Linear layer and the Sigmoid activation, which is applied to predict the GC representations for the multi-horizons.
III-G Loss Function
Thanks to the nature of the GC representations, the proposed FlightBERT++ framework is formulated as the MBC paradigm. In this work, the BCE loss function is employed to optimize the network parameters in the training procedure. Specifically, let the $p,y$ are the ground truth and prediction of a trajectory point, $p_{a_{j}}^{b}$, $y_{a_{j}}^{b}$ are the GC representation of the $j^{th},j\in[1,6]$, attribute in the ground truth and prediction, respectively. The calculation of the loss can be mathematically described as Eq. (9) - Eq. (10):
$\mathcal{L}(p,y)=\frac{1}{N}\sum_{n=1}^{N}\sum_{j=1}^{J}\mathcal{L}_{BCE}^{a_{j}}(p_{a_{j}}^{b},y_{a_{j}}^{b})$ (9)
$\mathcal{L}_{BCE}^{a_{j}}(p_{a_{j}}^{b},y_{a_{j}}^{b})=-\frac{1}{M}\sum_{i=1}^{M}(p_{a_{j}}^{b}[i]*log(y_{a_{j}}^{b}[i])+$ (10)
$(1-p_{a_{j}}^{b}[i])*log(1-y_{a_{j}}^{b}[i]))$
where the $N$ is the number of samples in the minibatch, while $M$ denotes the dimension of the GC representation of each attribute.
IV Experimental Settings
IV-A Dataset Preprocessing and Description
@@
To validate the proposed framework, a real-world flight trajectory dataset was collected from an ATC system in China [46]. The dataset contains a total of 9 days of trajectory data with 20 seconds intervals from February 19 to 27, 2021, in which the range of the interested (ROI) longitude and latitude are $[94.616,113.689]$ and $[19.305,37.275]$, respectively. The key attributes of the flight trajectory are extracted from the raw data to support the experiments, including timestamps, call sign, longitude, latitude, altitude, and velocity in $x,y,z$ directions. The positional components (longitude, latitude, and altitude) are in degree with the WGS-84 coordinate system, while the velocity in $x,y,z$ dimensions are measured by $km/h$ with the Cartesian coordinates.
After the aforementioned preprocess, a total of 8643 flight trajectories in the dataset and split into train, validate, and test subsets. Specifically, the trajectory of the first 7 days is used to train the FTP models while $8^{th}$ and $9^{th}$ are for validation and testing, respectively. The aforementioned data splits are applied to all the experiments in this work to ensure fair comparison.
IV-B Comparison Baselines
@@
In this work, a total of 7 competitive approaches serve as baselines to validate the effectiveness of the proposed FlightBERT++ framework. Moreover, the baselines are divided into two groups to conduct comparisons by multi-horizon prediction styles. Group A performs iterative multi-horizon prediction, i.e., the model only predicts the result for the next time step and serves as a pseudo-observation for multi-horizon inference. Group B is the direct multi-horizon prediction approach that forecasts the results in a single inference process for multiple time steps. The detailed descriptions of the baselines are listed as follows.
  • (A1) LSTM: An FTP framework using LSTM neural networks [38]. In this work, we employ a 4-layer LSTM network with 128 neurons to perform the FTP task.
  • (A2) Transformer: The Transformer block [12] server as backbone network to build the FTP model. The model is implemented by stacking 4 Transformer blocks with 128-dimensional hidden states and an FC layer.
  • (A3) Kalman-Filter: A typical model-driven flight state estimation algorithm based on historical observations and generally applied in target tracking scenarios. In this work, the Kalman-Filter (KF) is employed to perform the FTP task [22].
  • (A4) FlightBERT: A novel FTP framework proposed in our previous work, which converts the FTP tasks into MBC paradigm and achieves desired performance [9].
  • (A5) WTFTP: A time-frequency analysis FTP framework using wavelet transform, achieves higher precision and robustness, particularly in high-maneuvering flight scenarios. The model configuration is referred to the original WTFTP work, more details can be found in [11].
  • (B1) LSTM+Attention: A Seq2Seq architecture is designed to forecast the flight status across multiple future time steps directly. In this experiment, an Attention-based encoder-decoder LSTM network [47] is adapted to the FTP task. A 4-layer LSTM encoder, a 2-layer LSTM decoder, and a single-head attention module are employed to conduct the FTP model.
  • (B2) Transformer-Seq2Seq: In view of the Transformer achieving superior performance in sequential modeling tasks, a vanilla Seq2Seq Transformer architecture [12] is employed to implement the FTP task. In this model, the encoder and decoder are built by 4 and 2 Transformer blocks, with a 4-head multi-head attention mechanism, respectively.
IV-C Experimental Configuration
Based on ATC work, the quantization precision $0.001\degree$ (about 110 meters) is selected to adapt the transformation between GC representation and the decimals. In this resolution, the number of bits of the GC representation of trajectory attribute for inputs and differential values for outputs is listed as follows:
  • Longitude and latitude: we use 18 and 16 bits to encode the real values (decimals) of longitude and latitude into GC representation for the inputs. The number of the bits for longitude and latitude differential values are both set to 8 bits.
  • Altitude: the altitude is measured in $10m$, and encoded into the GC with 11 bits, while its differential value encodes to GC representation with 8 bits.
  • Velocity: the velocity is measured in Km/h and all are encoded into the GC representation with 11 bits for inputs. For the model outputs, the differential values of $Vx,Vy$ are represented by the 9-bit GC representation while the $Vz$ is encoded with a 6-bit GC representation.
Note that the highest bit of GC representations encoded from differential values indicates the sign of the real value. Based on the above configurations, the input of the proposed framework is a 78-dimensional vector while the output is a 48-dimensional vector. The embedding size of the trajectory attributes and horizons is set to 128. The number of the Transformer blocks in the trajectory encoder and differential-prompted decoder are set to 4 and 2, respectively. The number of hidden states of the Transformer blocks is set to 768. An attention operator with 4 heads is applied to all Transformer blocks in the proposed framework.
For experiments A1, A2, B1, and B2, the Z-Score normalization algorithm is applied to process the value into [0,1] for longitude and latitude attributes due to the sparse specificities of their distributions. While the altitude and velocities in $x,y,z$ dimension are normalized into [0,1] using the Max-min algorithm. The MSE loss function is applied to optimize the above models in the training process.
In the training phase, we use the latest 3-minute observations to predict the flight status of the future 5 minutes, i.e., predicting 15 trajectory points based on 9 observed trajectory points. The Adam optimizer with $10^{-4}$ initial learning rate is applied to train all the above deep learning-based models.
In this work, all the experiments are implemented with the open-source deep learning framework PyTorch 1.9.0. The models are trained on the server configured with Ubuntu 16.04 operating system, 8*NVIDIA GeForce RTX 2080 GPU, Intel(R) Core(TM) i7-7820X@3.6GHz CPU, and 128 GB memory.
IV-D Evaluation Metrics
In this work, the Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE) are applied to evaluate the proposed methods and baselines, which are the common criteria for the TP tasks. The definitions of the MAE, MPAE, and RMSE are described as follows:
$MAE=\frac{1}{N}\frac{1}{h}\sum_{i=1}^{N}\sum_{j=1}^{h}{\lvert a_{ij}-a^{\prime}_{ij}\rvert}$ (11)
$MAPE=\frac{1}{N}\frac{1}{h}\sum_{i=1}^{N}\sum_{j=1}^{h}{\lvert\frac{a_{ij}-a^{\prime}_{ij}}{a_{ij}}\rvert}\times 100\%$ (12)
$RMSE=\sqrt{\frac{1}{N}\frac{1}{h}\sum_{i=1}^{N}\sum_{j=1}^{h}\left(a_{ij}-a^{\prime}_{ij}\right)^{2}}$ (13)
where $N$ is the number of the samples in the test set, $h$ represents the prediction horizon. $a,a^{\prime}$ are the real-value (decimals) of trajectory attributes in the ground truth and prediction, respectively.
Although the longitude, latitude, and altitude (LLA) are measured using the above common metrics, it is not intuitive to evaluate the error separately due to the trajectory point being determined by LLA together in the three-dimensional (3D) airspace. To this end, the Mean Deviation Error (MDE) is proposed to evaluate the Euclidean distance ($km$) of the predictions and the actual trajectory points. Specifically, the LLA of both predictions and ground truth is projected to the earth-centered and earth-fixed (ECEF) coordinate system from the WGS-84 coordinate systems, and the deviation error in 3D airspace is measured to evaluate the model performance.
where $p,p^{\prime}$ are the transformed values of the ground truth and prediction in the ECEF coordinate system, respectively. $\Phi(\cdot)$ is the calculation function of Euclidean distance in the 3D airspace.
Furthermore, the Mean Time Costs (MTC) metric is proposed to evaluate the computational performance of the proposed framework and comparison baselines, defined as follows:
$MTC=\frac{1}{N}\sum_{i=1}^{N}time\_costs_{i}^{h}$ (15)
where $N$ is the number of samples in the testing process, $time\_costs_{i}^{h}$ represents the time cost for $h$ prediction horizons of sample $i$. In this phase, the batch size of all evaluation models is set to 1 to ensure comparison fairness.
TABLE I: The experimental results of the proposed framework and baselines.
Style Methods Horizon MAE $\downarrow$ MAPE (%) $\downarrow$ RMSE $\downarrow$ MDE $\downarrow$
@@ IV-D Evaluation Metrics
WTFTP (A5) 1 0.0021 0.0021 1.10 0.0020 0.0077 0.18 0.0049 0.0105 6.29 0.35
3 0.0034 0.0033 2.33 0.0032 0.0124 0.39 0.0077 0.0134 7.63 0.56
9 0.0109 0.0099 7.68 0.0102 0.0367 1.21 0.0233 0.0265 16.76 1.71
15 0.0211 0.0187 14.40 0.0197 0.0693 2.19 0.0429 0.0432 27.80 3.28
Direct LSTM+Attention (B1) 1 0.0064 0.0068 1.98 0.0059 0.0249 0.30 0.0109 0.0138 5.96 1.09
@@
FlightBERT++ (Proposed) 1 0.0012 0.0012 0.93 0.0011 0.0046 0.17 0.0031 0.0071 8.86 0.22
3 0.0023 0.0023 1.92 0.0022 0.0086 0.35 0.0059 0.0115 10.05 0.39
9 0.0058 0.0058 4.69 0.0054 0.0215 0.84 0.0156 0.0213 15.58 0.97
15 0.0091 0.0090 6.59 0.0085 0.0336 1.19 0.0239 0.0298 20.25 1.54
V Result and Discussions
V-A Results and Quantitative Analysis
V-A1 Overall Performance of FTP